OpenStudy (anonymous):

Does (x-3)/(2x^2-5) not equal -3/-5 when x=0? I realize the end result will be 0.6 no matter what but every thing I use to check this result gives me 3/5 instead of -3/-5.

OpenStudy (anonymous):

Your answer is correct and confusion genuine. Whenever the sign notations of the denominator and numerator are same they are not mentioned explicitly, but implied through their absence. the arithmetic that leads to this result is as follows \[\frac{-3\times (-1)}{-5\times (-1)} = \frac{+3}{+5}= 0.6\] similarly\[\frac{+3}{+5}= \frac{3}{5}= 0.6\]

OpenStudy (anonymous):

we never write\[\frac{3}{5}\] as \[\frac{+3}{+5}\] the similarity of the signs is implied, it is understood that the signs could also have been - each and simply cancelled out as shown above.

OpenStudy (anonymous):

Thank you very much. I was worried that I had been missing some basic aspect of my math because of my tiredness, so I thought I should run it by someone else to check. Thanks!

OpenStudy (unklerhaukus):

-3/-5 = 3/5

OpenStudy (anonymous):

Also, thank you for the in-depth explanation as well.

OpenStudy (anonymous):

@Jaeriko You are welcome :)